A wheel has a radius of 3.90 m. How far (path length) does a point on the circumference travel if the wheel is rotated through the following angles, respectively? (a) 28.0° m (b) 28.0 rad m (c) 28.0 rev
Answer #1

Distance traveled by a point on circumference will be:

S = r*

Where r = radius of wheel = 3.90 m

= Angle in radian

Part A.

when = 28.0 degree

= 28.0*(pi/180) rad = 0.4887 rad

So,

S = 3.90*0.4887 = 1.91 m

Part B.

when = 28.0 rad

then

S = 3.90*28.0 = 109.20 m

Part C.

when = 28.0 rev

= 28.0 rev*(2*pi rad/1 rev) = 175.93 rad

So,

S = 3.90*175.93

S = 686.13 m

Please Upvote.